OpenStudy (anonymous):

solution find x and y y-x=-13 -4x+3y=-51

OpenStudy (hitaro9):

Alright, there are two ways you can look at solving a problem like this. You can either go with substitution, or you can go with equation addition. Have you gone over both? If so, is either more intuitive for you?

OpenStudy (anonymous):

oh yeah, my bad its substitution

OpenStudy (hitaro9):

Alright. Equation addition works just as well, but if you haven't gone over that yet then I wouldn't want to confuse you.

OpenStudy (hitaro9):

Alright, so the basic principle behind substitution is that you're trying to eliminate a letter by substituting in part of the equation from another part. So we look at this and we see the first equation y-x=-13 can be very easily manipulated to y=-13+x

OpenStudy (hitaro9):

Now this is certainly known to be true. This is a fact and we can take this equality and use it anywhere. Anyplace we see "-13 +x" we can put a y, and anywhere we see a "y" we can put in -13+x

OpenStudy (hitaro9):

So we look at our second equation -4x + 3y =-51 and see we can entirely eliminate y by substituting in -13+x wherever it is.

OpenStudy (hitaro9):

So if we go ahead and try that we have -4x+3(-13+x) =-51

OpenStudy (hitaro9):

And from there, it's simple algebra. Do you think you can go through that on your own?

OpenStudy (anonymous):

im not sure, would the first step end up being -4x+3-13x=-51?

OpenStudy (hitaro9):

Umm no. Remember you can always go back to order of operations if need be.

OpenStudy (hitaro9):

Umm, you're familiar with 3(-13+x) notation right? if you see a number outside that parenthesis that means multiplication

OpenStudy (anonymous):

so i multiply 3*-13

OpenStudy (hitaro9):

So why don't you show me, if the problem were just -4x+3(-13+x) = -51 How would you go about finding x?

OpenStudy (hitaro9):

If you're multiplying two things, you've got to remember to distribute. Are you familiar with how to work with parenthesis? Or have you not seen something written like that before?

OpenStudy (anonymous):

i have , but a few years ago. im in geometry but i have to get a credit for algebra. and so i forgot how it all really works

OpenStudy (hitaro9):

Alright that's fair. If having letters thrown in there makes it seem more complicated, you can always take it back to something simpler. That's something I like to do if I'm ever lost. Like, if you're ever confused about fractions or multiplication rules, just use really simple examples.

OpenStudy (hitaro9):

So say x is 2 and we want to know what 3(-13+2) is.

OpenStudy (anonymous):

-39+6?

OpenStudy (hitaro9):

If we just say it's -13(3) +2 we end up with -39 +2 = -37 if we multiply everything in the parenthesis we have -13(3) +2(3) = -39 +6 = -33

OpenStudy (hitaro9):

Right. So you can just double check if you're not sure about multiplication rules

OpenStudy (hitaro9):

So you want to multiply everything in the parentthesis by 3

OpenStudy (hitaro9):

So 3(-13+x) = -13(3) +x(3)

OpenStudy (hitaro9):

=-39 +3x

OpenStudy (hitaro9):

-4x+3(-13+x) =-51 -4x -39 +3x =-51

OpenStudy (hitaro9):

Do you think you can solve it from there?

OpenStudy (anonymous):

most likely , but since there are two of the x will i need to add 39 to the other side

OpenStudy (hitaro9):

Right. So adding 39 to both sides we have .... -4x -39 +3x +39 = -51 +39

OpenStudy (hitaro9):

Which you can simplify to -4x +3x = -12

OpenStudy (anonymous):

and then i add -4x and 3x and then divid it by 12?

OpenStudy (hitaro9):

Let's take it 1 step at a time. First add the -4x and 3x

OpenStudy (hitaro9):

What should the equation look like now?

OpenStudy (anonymous):

-1x=12

OpenStudy (hitaro9):

That's right. Remember, we are trying to find what x is. Can you see what x is from that?

OpenStudy (anonymous):

-12

OpenStudy (hitaro9):

Oh wait. Close. Small mistake you dropped a negative sign from the -12 earlier.

OpenStudy (hitaro9):

Scroll up and you'll see the equation was -4x +3x = -12

OpenStudy (hitaro9):

So the answer is going to be x =12 right?

OpenStudy (anonymous):

oh okay so 12

OpenStudy (hitaro9):

Alright. So we have half the answer, and the next step is super simple. Let's rewrite the problem: y-x=-13 -4x+3y=-51 and we know that x = 12

OpenStudy (hitaro9):

Do you see how you can (fairly quickly) find what y is from that?

OpenStudy (anonymous):

would it be -1

OpenStudy (hitaro9):

Right. Perfect.

OpenStudy (hitaro9):

So then you have your answers y = -1 x= 12

OpenStudy (hitaro9):

Just to be sure we can plug these in to each equation to make sure we have the right answer.

OpenStudy (hitaro9):

-1)-(12)=-13 -1-12 =-13 -13 =-13 which is certainly true -4(12)+3(-1) = -51 -48 -3 =-51 -51 = -51 Which is certainly true

OpenStudy (hitaro9):

Do you understand all the steps we took to get to that, or do you want to try another one?

OpenStudy (anonymous):

i get it now :)

OpenStudy (hitaro9):

Alright cool c: That took a while. If you wanted to show your work concisely it would look something like -4x+3y=-51 y-x =-13 y= -13+x Substituting -> -4x+3(-13+x) =-51 -4x -39 +3x = -51 -x =-12 x=12 Subtituting -> y-(12) =-13 y=-1 Which only took about a minute or two to type

OpenStudy (anonymous):

alright, thank you though